OpenStudy (anonymous):

Solve the triangle. A = 33°, a = 19, b = 14

OpenStudy (jdoe0001):

I assume you've covered the "law of sines" already thus 

OpenStudy (anonymous):

i understand how to set everything up 19/sin33=14/sinB 19sinB=14sin33 sinB=14sin33/19 sinB=.4013 now what i dont understand is how to get it in degrees

OpenStudy (jdoe0001):

\(\bf \cfrac{sin(A)}{a}=\cfrac{sin(B)}{b}\implies \cfrac{sin(33^o)}{19}=\cfrac{sin(B)}{14} \\ \quad \\ \cfrac{sin(33^o)\cdot 14}{19}=sin(B)\implies sin^{-1}\left[\cfrac{sin(33^o)\cdot 14}{19}x\right]=sin^{-1}[sin(B)] \\ \quad \\ sin^{-1}\left[\cfrac{sin(33^o)\cdot 14}{19}x\right]=\measuredangle B\)

OpenStudy (jdoe0001):

hmmm I have an "x" hitchhiking there as a typo darn anyhow \(\bf \cfrac{sin(A)}{a}=\cfrac{sin(B)}{b}\implies \cfrac{sin(33^o)}{19}=\cfrac{sin(B)}{14} \\ \quad \\ \cfrac{sin(33^o)\cdot 14}{19}=sin(B)\implies sin^{-1}\left[\cfrac{sin(33^o)\cdot 14}{19}\right]=sin^{-1}[sin(B)] \\ \quad \\ sin^{-1}\left[\cfrac{sin(33^o)\cdot 14}{19}\right]=\measuredangle B\)

OpenStudy (jdoe0001):

notice that your 33 are in degrees, thus make sure your calculator is in Degree mode

OpenStudy (anonymous):

thank you now i understand how to get the value of angle B

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

yes i have my mode set in degrees